# ExoMiner++ 2.0: Vetting TESS Full-Frame Image Transit Signals

### v1

@@ V.1 Experimental Setup
To evaluate the model under different data splits, we conducted 5-fold cross-validation (CV) experiments, providing a more reliable estimate of performance across multiple splits. This approach ensures that all available data are used for both training and evaluation and helps assess generalization more robustly. The data are split into five folds such that the TCEs associated with the same target are included in the same fold. Given that all TCEs generated from the same target light curve and pixel data are either part of the training or test sets in any given CV iteration, we minimize data leakage. Furthermore, to ensure that the folds have approximately the same distribution of planets and non-planets as the overall dataset, we use a greedy strategy to assign TCEs to cross-validation folds. For each target, all its TCEs are grouped together and assigned to the fold that holds the fewest planet TCEs. This iterative process ensures that the distribution of planet across folds remains as balanced as possible.
@@
To formally assess whether one model outperforms the other, we use a paired, stratified, instance-level bootstrap on the combined out-of-fold predictions from all five test sets. This approach resamples instances with class-stratification and computes the difference in each metric ($\Delta=metric_{ModelB}-metric_{ModelA}$) for each bootstrap replicate. From these replicates, we derive percentile confidence intervals (CIs) for the difference and assess both statistical significance (i.e., whether the CI excludes zero) and practical significance (i.e., whether CI clears a pre-specified effect size threshold $\epsilon$. This method leverages all available out-of-fold predictions rather than reducing the comparison to five fold-level aggregates, leading to substantially greater statistical efficiency in large-sample settings relative to fold-level tests. It avoids parametric assumptions about the distribution of fold means and exploits the paired nature of per-instance predictions, while conditioning on the fitted models obtained from the CV experiments (Efron & Tibshirani, 1994).
We computed 5,000 stratified bootstrap replicates, and estimated the 95% confidence intervals for each metric (i.e., significance level set to $\alpha=0.05$). The effect size thresholds were chosen to reflect operational impact (e.g., 1% increase in performance). For all proportion-based metrics (i.e., recall, precision, and accuracy) and PR AUC, the effect size threshold was set to $\epsilon=0.01$ , which means that a 1% difference is considered practically significant. As for ROC AUC, the threshold was set to $\epsilon=0.005$ since it is less sensitive to class imbalance.
For each cross-validation iteration, we train a set of 10 models and construct an ensemble by averaging their prediction scores using the mean. This approach helps reduce variability caused by random weight initialization and the inherent stochasticity of mini-batch gradient descent during optimization. By training multiple models with different weight initializations, the ensemble captures variations in local minima encountered during the optimization of the loss function, resulting in a more accurate estimate of the model’s performance.
All models are trained for 300 epochs, optimized using binary cross-entropy as the loss function and the Adam optimizer (Kingma, 2014) ($\beta_{1}=0.900$, $\beta_{2}=0.999$, $\epsilon=1\mathrm{e}{-8}$). Model selection is performed using early stopping with a patience of 20 epochs to prevent overfitting; i.e., if no improvement in the validation PR AUC is observed after 20 epochs, the training is stopped, and the model instance at that point is selected as the final model. The CV dataset can be found on Zenodo (Martinho & Valizadegan, 2025).
@@ V.2 Multi-source Learning
As the ongoing SPOC planet searches produce results for more sector runs, the number of FFI TCEs is expected to grow and surpass the 2-minute results. However, the current number of sector runs performed for the FFI data is significantly smaller than those for the 2-minute data, and only a single multi-sector run has been conducted as mentioned in Section III.2. For this reason, using the 2-minute labeled dataset for training the model might result in improved performance on FFI data. Following the work in Valizadegan et al. (2025), we conducted multi-source learning (Zhuang et al., 2020) using as sources the TESS SPOC 2-min and FFI TCEs datasets for training.
The FFI and 2-min datasets differ in a few key important ways: 1) Cadence: the 2-minute data have a shorter cadence, whereas TESS full-frame images are sampled at either 30min30\ $\text${min}, 10min10\ $\text${min}, or 200s200\ $\text${s} depending on the sector. The SPOC pipeline resamples the light curves to a common cadence of 10 minutes for the multi-sector runs since the S14–S55 search, partially reducing the difference; 2) Systematics correction: the 2-minute light curves benefit from more effective systematics correction, including less blending and crowding on average; 3) Vetting quality: the 2-minute data have higher-quality dispositions due to more extensive and focused vetting of transit signals. Although the 2-min targets also appear in FFI data, the shorter and fewer sector runs for these data limit the detection of TCEs by the SPOC pipeline, particularly for longer-period TOIs. Additionally, targets observed exclusively in FFI data have received less scrutiny to date; and 4) Target population: FFI data cover a more diverse target population, including those targets already observed for 2-minute data as well as many more fainter targets.
Multi-source learning allows the model to simultaneously learn shared representations in a larger combined dataset that generalizes across both domains (i.e., the two collection modes), which can be more effective than fine-tuning (e.g., pre-training on the 2-min dataset and then continuing training on the FFI dataset) since that may overwrite useful features learned from the 2-min data. In this setting, the model is exposed during training to different noise levels and target populations which provides better generalization to new, unseen sectors.
@@ V.3 Results and Discussion
V.3.1 Cross-Validation Results
Table 3 presents the overall performance of the ExoMiner++ and ExoMiner++ 2.0 models across a multi-source 5-fold cross-validation procedure, evaluated using standard binary classification metrics. The first thing to notice is that the results for ExoMiner++ are in line with those reported in Valizadegan et al. (2025) when the model is trained on TESS 2-min data. Even though the dataset used in the current work has a different data distribution and includes more sector runs and FFI TCEs, the overall performance metrics remain consistent.
@@ V.3.1 Cross-Validation Results
Since we aim to assess whether incorporating neighbors’ data improves the model’s ability to reduce confidence for TCEs with observable transit-source offsets near bright stars, we also examined the recall of both models for nearby false positives, specifically TOIs classified as nearby planet candidates and eclipsing binaries in SG1. The recall for these nearby FPs is 0.92 and 0.94 for ExoMiner++ and ExoMiner++ 2.0, respectively. Figure 5 shows the distribution of scores for the two models for both planet and nearby FP TCEs. The shift in both distributions indicates that ExoMiner++ 2.0 lowered the scores of all objects, true planets or background transits. As such, it is more conservative.
@@
A similar behavior is also observed when we plot the mean average score for planets and nearby FPs as a function of the CROWDSAP metric (Figure 6). TCEs observed in multiple sectors were excluded from this analysis because, in those cases, there are multiple CROWDSAP metrics available, one for each sector where the target was observed, and there is no straightforward method to aggregate these metric values into a single statistic. The CROWDSAP metric in TESS data quantifies the fraction of flux in the photometric aperture that originates from the target star; values near unity indicate minimal contamination, whereas lower values indicate significant flux contribution from nearby stars or background sources. As the contamination increases (i.e., lower CROWDSAP values), both models exhibit a decrease in the planet average score, while the average score for nearby FPs remains consistent (results below CROWDSAP $\approx 0.5$ are less reliable given that those bins contain fewer than 100 TCEs). However, even in cases where nearly all of the flux in the aperture comes from the target star (i.e., CROWDSAP close to 1), ExoMiner++ 2.0 still shows an average lower score for both planet and nearby FP TCEs. This conservativeness translates into a tradeoff, as the model degrades its performance on planet TCEs to correctly classify nearby false positives. Three hypotheses can explain this: 1) the large pixel scale (21arcsec21\ arcsec) means that even in moderately crowded fields, the light from the target star is often mixed with nearby stars that fall into the same photometric aperture or even the same pixel and can still lead to errors in the transit source location estimation even when conducting difference image analysis. This limitation constitutes a lower bound that ultimately limits the capability of the model to separate on-target transits from background sources; 2) the small number of nearby FPs relative to the size of the dataset, combined with the joint optimization of the model means that it is challenging for the model to optimize its ‘Difference Image’ branch to learn features useful for detecting real transit source offsets. For example, pretraining the difference image branch on a dedicated task that distinguishes on-target transit signals from those originating in nearby stars could enable the branch to learn discriminative representations for this scenario; 3) no dedicated hyperparameter optimization (HPO) was performed for this architecture, particularly the ‘Difference Image’ branch. We expect that by conducting a new HPO run we can find a more optimal architecture that can use the difference image data more effectively, including the neighbors information. Other potential future improvements to enhance the performance of the difference image branch in cases where the transit source may be a neighboring star include incorporating a lightweight spatial attention mechanism inspired by (Woo et al., 2018) to guide the model toward salient regions of the image such as the transit signal location, while suppressing irrelevant areas like pixels without known stars. Another promising direction is to select difference images from sectors with higher quality metrics rather than relying on random sampling with replacement, ensuring the model is exposed to a set of higher-quality inputs.
@@ V.3.2 Comparison Between Data Sources
As outlined in Section V.2, the 2-min and FFI datasets exhibit distinct observational and noise characteristics, which can influence model behavior. In this section, we start by comparing the overall performance on these two datasets. To remove confounding effects from this analysis, we exclude NTP TCEs since this subset is dependent on the availability of TEC’s flux triage results. As mentioned in Section III.3, we had access to an inconsistent set of these results: for 2-min data we used the flux triage results for single-sector runs S1 through S41, while for FFI we had results for single-sector runs S40 through S73. Besides not having the results for the same sector runs, the numbers of NTP TCEs generated by these two results varies significantly as shown in Tables 1 and 2. Furthermore, the SPOC planet searches for FFI data start only with Sector 36, and so far only one multi-sector run was conducted. Because of this, the number of multi-sector runs from 2-min data is significantly larger, and on average longer than the single multi-sector run conducted for FFI data (S56–S69). With this bias in mind, for this analysis we considered only TCEs from the planet searches that use the same sectors, which in this case are all single-sector runs from Sector 36 through Sector 72, totaling 37 runs. Given that the ephemeris matching was conducted for the same sector runs and using the same label sources, the only differences in these two sets come from the target selection criteria for FFI and 2-min data, and the operation of the SPOC pipeline when performing planet searches in these two data collection modes. Table 5 shows the number of TCEs, targets, and events (i.e., TOIs and Prsa’s EBs) for 2-min and FFI data after applying these filters.
Table 5: Comparison between 2-min and FFI datasets. NTP TCEs were excluded and only TCEs from the same sector runs were considered from both datasets.
Category 2-min FFI
@@
As expected, the FFI data contain more TCEs as the number of targets selected for FFI planet searches was larger than for 2-min. Following this trend, the number of events matched to TCEs was also larger. However, both datasets have targets and events that are exclusive to each other. In the case of events, the number of these objects that are shared between 2-min and FFI datasets is 1,708, while 533 and 388 are found exclusively in FFI and 2-min data, respectively. Interestingly, the number of detected KP TCEs in 2-min data is significantly larger than for FFI data, while a similar trend is shown for CP TCEs but favoring FFI data. Besides those differences, the FFI dataset also shows significantly more FP TCEs detected than the 2-min counterpart.
Table 6: Binary classification performance for ExoMiner++ 2.0 from CV experiment computed separately for the 2-min and FFI datasets. Precision, recall, and accuracy are computed at a classification threshold of 0.5. The best performer is highlighted in bold. Point estimates are computed across the CV iterations by aggregating all test sets.
Source Precision & Recall PR AUCaaThese metrics should be compared with caution because they are sensitive to class distribution differences between datasets. Normalized PR AUC ROC AUC AccuracyaaThese metrics should be compared with caution because they are sensitive to class distribution differences between datasets. Precision @ 95% RecallaaThese metrics should be compared with caution because they are sensitive to class distribution differences between datasets. Recall @ 95% PrecisionaaThese metrics should be compared with caution because they are sensitive to class distribution differences between datasets.
2-min 0.937 & 0.845 0.967 0.957 0.989 0.948 0.876 0.767
@@
Table 6 shows the performance metrics for ExoMiner++ 2.0 on these two datasets. The metrics were computed by aggregating the test set predictions across all CV iterations. Overall, the 2-min dataset shows higher performance, including on normalized PR AUC and ROC AUC, both of which are considered less sensitive to variations in class distribution across the two datasets. The normalized PR AUC is computed as:
@@
where $\gamma$ denotes the prevalence of positive examples in the dataset.
To investigate the causes behind the lower performance in the FFI dataset, we examined parameters that could explain these differences. More concretely, we analyzed the distribution of targets as function of $T_{mag}$ and Gaia RUWE, as well as the distribution of TCEs as function of their multiple event statistic (MES) and TCE model SNR. None of these parameters showed a clear shift in distribution between the two datasets. However, the distribution of the CROWDSAP metric revealed a noticeable shift for FFI targets toward lower values (Figure 7). A substantial fraction of FFI targets exhibit CROWDSAP values below 0.9, meaning that at least 10% of the flux in the aperture originates from other sources. Such contamination can dilute transit depths, bias planet radius estimates, reduce MES and SNR, and introduce systematics that are more difficult to correct by algorithms such as the PDC module in the SPOC pipeline. Furthermore, blended eclipsing binaries and variable stars within the aperture can mimic transit-like signals, increasing the false positive rate. Although MES and SNR distributions do not appear strongly affected, the deterioration in ExoMiner++ 2.0 performance suggests that contamination impacts other classification-relevant features, such as transit shape and detrending quality.
@@
Figure 7: Target population for the 2-min and FFI TCE datasets as a function of CROWDSAP Metric.

### v2

@@ V.1 Experimental Setup
To evaluate the model under different data splits, we conducted 5-fold cross validation (CV) experiments, providing a more reliable estimate of performance across multiple splits. This approach ensures that all available data are used for both training and evaluation and helps assess generalization more robustly. The data are split into five folds such that the TCEs associated with the same target are included in the same fold. Given that all TCEs generated from the same target light curve and pixel data are either part of the training or test sets in any given CV iteration, we minimize data leakage. Furthermore, to ensure that the folds have approximately the same distribution of planets and non-planets as the overall dataset, we use a greedy strategy to assign TCEs to cross validation folds. For each target, all its TCEs are grouped together and assigned to the fold that holds the fewest planet TCEs. This iterative process ensures that the distribution of planet across folds remains as balanced as possible.
@@
For each cross validation iteration, we train a set of 10 models and construct an ensemble by averaging their prediction scores using the mean. This approach helps reduce variability caused by random weight initialization and the inherent stochasticity of mini-batch gradient descent during optimization. By training multiple models with different weight initializations, the ensemble captures variations in local minima encountered during the optimization of the loss function, resulting in a more accurate estimate of the model’s performance.
All models are trained for 300 epochs, optimized using binary cross-entropy as the loss function and the Adam optimizer (adam2014method) ($\beta_{1}=0.900$, $\beta_{2}=0.999$, $\epsilon=1\mathrm{e}{-8}$). Model selection is performed using early stopping with a patience of 20 epochs to prevent overfitting; i.e., if no improvement in the validation PR AUC is observed after 20 epochs, the training is stopped, and the model instance at that point is selected as the final model. The CV dataset can be found on Zenodo (Martinho & Valizadegan, 2025).
@@ V.2 Multi-source Learning
As the ongoing SPOC planet searches produce results for more sector runs, the number of FFI TCEs is expected to grow and surpass the 2-minute results. However, the current number of sector runs performed for the FFI data is significantly smaller than those for the 2-minute data, and only a single multisector run has been conducted as mentioned in Section III.2. For this reason, using the 2-minute labeled dataset for training the model might result in improved performance on FFI data. Following the work in Valizadegan et al. (2025), we conducted multi-source learning (Zhuang et al., 2020) using as sources the TESS SPOC 2-minute and FFI TCEs datasets for training.
The FFI and 2-minute datasets differ in a few key important ways: 1) Cadence: the 2-minute data have a shorter cadence, whereas TESS full-frame images are sampled at either 30min30\ $\text${min}, 10min10\ $\text${min}, or 200s200\ $\text${s} depending on the sector. The SPOC pipeline resamples the light curves to a common cadence of 10 minutes for the multisector runs since the S14–S55 search, partially reducing the difference; 2) Systematics correction: the 2-minute light curves benefit from more effective systematics correction, including less blending and crowding on average; 3) Vetting quality: the 2-minute data have higher-quality dispositions due to more extensive and focused vetting of transit signals. Although the 2-minute targets also appear in FFI data, the shorter and fewer sector runs for these data limit the detection of TCEs by the SPOC pipeline, particularly for longer-period TOIs. Additionally, targets observed exclusively in FFI data have received less scrutiny to date; and 4) Target population: FFI data cover a more diverse target population, including those targets already observed for 2-minute data as well as many more fainter targets.
Multi-source learning allows the model to simultaneously learn shared representations in a larger combined dataset that generalizes across both domains (i.e., the two collection modes), which can be more effective than fine-tuning (e.g., pre-training on the 2-minute dataset and then continuing training on the FFI dataset) since that may overwrite useful features learned from the 2-minute data. In this setting, the model is exposed during training to different noise levels and target populations which provides better generalization to new, unseen sectors.
@@ V.3 Results and Discussion
V.3.1 Cross Validation Results
Table 3 presents the overall performance of the ExoMiner++ and ExoMiner++ 2.0 models across a multi-source 5-fold cross validation procedure, evaluated using standard binary classification metrics. The first thing to notice is that the results for ExoMiner++ are in line with those reported in Valizadegan et al. (2025) when the model is trained on TESS 2-minute data. Even though the dataset used in the current work has a different data distribution and includes more sector runs and FFI TCEs, the overall performance metrics remain consistent.
@@ V.3.1 Cross Validation Results
Since we aim to assess whether incorporating neighbors’ data improves the model’s ability to reduce confidence for TCEs with observable transit-source offsets near bright stars, we also examined the recall of both models for nearby false positives, specifically TOIs classified as nearby planet candidates and eclipsing binaries in SG1. The recall for these nearby FPs is 0.92 and 0.94 for ExoMiner++ and ExoMiner++ 2.0, respectively. Figure 5 shows the distribution of scores for the two models for both planet and nearby FP TCEs. The shift in both distributions indicates that ExoMiner++ 2.0 lowered the scores of all objects, true planets or background transits, thus confirming that this model is more conservative than ExoMiner++.
@@
A similar behavior is also observed when we plot the mean average score for planets and nearby FPs as a function of the CROWDSAP metric (Figure 6). TCEs observed in multiple sectors were excluded from this analysis because, in those cases, there are multiple CROWDSAP metrics available, one for each sector where the target was observed, and there is no straightforward method to aggregate these metric values into a single statistic. The CROWDSAP metric in TESS data quantifies the fraction of flux in the photometric aperture that originates from the target star; values near unity indicate minimal contamination, whereas lower values indicate significant flux contribution from nearby stars or background sources. As the contamination increases (i.e., lower CROWDSAP values), both models exhibit a decrease in the planet average score, while the average score for nearby FPs remains consistent (results below CROWDSAP $\approx 0.5$ are less reliable given that those bins contain fewer than 100 TCEs). However, even in cases where nearly all of the flux in the aperture comes from the target star (i.e., CROWDSAP close to 1), ExoMiner++ 2.0 still shows an average lower score for both planet and nearby FP TCEs. This conservativeness translates into a tradeoff, as the model degrades its performance on planet TCEs to correctly classify nearby false positives. Three hypotheses can explain this: 1) the large pixel scale ($21\arcsec$) means that even in moderately crowded fields, the light from the target star is often mixed with nearby stars that fall into the same photometric aperture or even the same pixel and can still lead to errors in the transit source location estimation even when conducting difference image analysis. This limitation constitutes a lower bound that ultimately limits the capability of the model to separate on-target transits from background sources; 2) the small number of nearby FPs relative to the size of the dataset, combined with the joint optimization of the model means that it is challenging for the model to optimize its ‘Difference Image’ branch to learn features useful for detecting real transit source offsets. For example, pretraining the ‘Difference Image’ branch on a dedicated task that distinguishes on-target transit signals from those originating in nearby stars could enable the branch to learn discriminative representations for this scenario; 3) no dedicated hyperparameter optimization (HPO) was performed for this architecture, particularly the ‘Difference Image’ branch. We expect that by conducting a new HPO run we can find a more optimal architecture that can use the difference image data more effectively, including the neighbors information. Other potential future improvements to enhance the performance of the ‘Difference Image’ branch in cases where the transit source may be a neighboring star include incorporating a lightweight spatial attention mechanism inspired by (Woo et al., 2018) to guide the model toward salient regions of the image such as the transit signal location, while suppressing irrelevant areas like pixels without known stars. Another promising direction is to select difference images from sectors with higher quality metrics rather than relying on random sampling with replacement, ensuring the model is exposed to a set of higher-quality inputs.
@@ V.3.2 Comparison Between Data Sources
As outlined in Section V.2, the 2-minute and FFI datasets exhibit distinct observational and noise characteristics, which can influence model behavior. In this section, we start by comparing the overall performance on these two datasets. To remove confounding effects from this analysis, we exclude NTP TCEs since this subset is dependent on the availability of TEC’s flux triage results. As mentioned in Section III.3, we had access to an inconsistent set of these results: for 2-minute data we used the flux triage results for single-sector runs S1 through S41, while for FFI we had results for single-sector runs S40 through S73. Besides not having the results for the same sector runs, the numbers of NTP TCEs generated by these two results varies significantly as shown in Tables 1 and 2. Furthermore, the SPOC planet searches for FFI data start only with Sector 36, and so far only one multisector run was conducted. Because of this, the number of multisector runs from 2-minute data is significantly larger, and on average longer than the single multisector run conducted for FFI data (S56–S69). With this bias in mind, for this analysis we considered only TCEs from the planet searches that use the same sectors, which in this case are all single-sector runs from Sector 36 through Sector 72, totaling 37 runs. Given that the ephemeris matching was conducted for the same sector runs and using the same label sources, the only differences in these two sets come from the target selection criteria for FFI and 2-minute data, and the operation of the SPOC pipeline when performing planet searches in these two data collection modes. Table 5 shows the number of TCEs, targets, and events (i.e., TOIs and Prsa’s EBs) for 2-minute and FFI data after applying these filters.
Table 5: Comparison between 2-minute and FFI datasets. NTP TCEs were excluded and only TCEs from the same sector runs were considered from both datasets.
Category 2-minute FFI
@@
As expected, the FFI data contain more TCEs as the number of targets selected for FFI planet searches was larger than for 2-min. Following this trend, the number of events matched to TCEs was also larger. However, both datasets have targets and events that are exclusive to each other. In the case of events, the number of these objects that are shared between 2-minute and FFI datasets is 1,708, while 533 and 388 are found exclusively in FFI and 2-minute data, respectively. Interestingly, the number of detected KP TCEs in 2-minute data is significantly larger than for FFI data, while a similar trend is shown for CP TCEs but favoring FFI data. Besides those differences, the FFI dataset also shows significantly more FP TCEs detected than the 2-minute counterpart.
Table 6: Binary classification performance for ExoMiner++ 2.0 from CV experiment computed separately for the 2-minute and FFI datasets. Precision, recall, and accuracy are computed at a classification threshold of 0.5. The best performer is highlighted in bold. Point estimates are computed across the CV iterations by aggregating all test sets.
Source Precision & RecallaaThese metrics should be compared with caution because they are sensitive to class distribution differences between datasets. PR AUCaaThese metrics should be compared with caution because they are sensitive to class distribution differences between datasets. Normalized PR AUC ROC AUC AccuracyaaThese metrics should be compared with caution because they are sensitive to class distribution differences between datasets. Precision @ 95% RecallaaThese metrics should be compared with caution because they are sensitive to class distribution differences between datasets. Recall @ 95% PrecisionaaThese metrics should be compared with caution because they are sensitive to class distribution differences between datasets.
2-minute 0.937 & 0.845 0.967 0.957 0.989 0.948 0.876 0.767
@@
Table 6 shows the performance metrics for ExoMiner++ 2.0 on these two datasets. The metrics were computed by aggregating the test set predictions across all CV iterations. Overall, the 2-minute dataset shows higher performance, including on normalized PR AUC and ROC AUC, both of which are considered less sensitive to variations in class distribution across the two datasets. The normalized PR AUC is computed as:
@@
where $\gamma$ denotes the prevalence of positive examples in the dataset. This normalization maps the random classifier baseline to zero and a perfect classifier to 1.
To investigate the causes behind the lower performance in the FFI dataset, we examined parameters that could explain these differences. More concretely, we analyzed the distribution of targets as function of $T_{mag}$ and Gaia RUWE, as well as the distribution of TCEs as function of their multiple event statistic (MES) and TCE model SNR. None of these parameters showed a clear shift in distribution between the two datasets. However, the distribution of the CROWDSAP metric revealed a noticeable shift for FFI targets toward lower values (Figure 7). A non-negligible fraction of FFI targets exhibit CROWDSAP values below 0.9, meaning that at least 10% of the flux in the aperture originates from other sources. Such contamination can dilute transit depths, bias planet radius estimates, reduce MES and SNR, and introduce systematics that are more difficult to correct by algorithms such as the PDC module in the SPOC pipeline. Furthermore, blended eclipsing binaries and variable stars within the aperture can mimic transit-like signals, increasing the false positive rate. Although MES and SNR distributions do not appear strongly affected, the deterioration in ExoMiner++ 2.0 performance suggests that contamination impacts other classification-relevant features, such as detrending quality.
@@
Figure 7: Target population for the 2-minute and FFI TCE datasets as a function of CROWDSAP Metric.